A five-year project has an initial fixed asset investment of $320,000, an initial NWC investment of $32,000, and an annual OCF of −$31,000. The fixed asset is fully depreciated over the life of the project and has no salvage value. If the required return is 10 percent, what is this project’s equivalent annual cost, or EAC? (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Equivalent annual cost $   
Answer #1

Solution:

Calculation of the Equivalent Annual Cost (EAC):

To calculate the EAC of the project, we first need the NPV of the project.

Notice that we include the NWC expenditure at the beginning of the project, and recover the NWC at the end of the project. The NPV of the project is:

NPV=-\$320,000-\$32,000-\$31,000(PVIFA_{10\%,5})+\frac{\$32,000}{(1.10)^{5}}

NPV=-\$320,000-\$32,000-\$31,000(3.7908)+\frac{\$32,000}{1.61051}

NPV=-\$320,000-\$32,000-\$117,514.8+\$19,869.48

NPV=-\$449,645.32

Now we can find the EAC of the project. The EAC is:

Equivalent Annual Cost (EAC)= \frac{-449,645.32}{PVIFA_{10\%,5}}

Equivalent Annual Cost (EAC)= \frac{-449,645.32}{3.7908}

Equivalent Annual Cost (EAC)= -\$118,614.89

Therefore, the Equivalent Annual Cost (EAC) is -$118,614.89